OpenStudy (anonymous):

(-3,-9) (1,-1) calculate the gradient of the line

OpenStudy (kira_yamato):

Gradient = (-1 - (-9)) / (1 - (-3)) Gradient = 8/4 =2 Gradient = 2

OpenStudy (luffingsails):

Gradient is another word for slope. So, just calculate the slope: \[\frac{rise}{run} = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]